If the roots of a1x2 +b1x +c1=0 are m1,m2  and the roots of  a2x2 +b2x +c2=0   are n1,n2   such that    m1 n1 = m2n2  =1, then


 


a)    a1/ a2= b1/ b2= c1/ c2


b)    a1 a2= b1 b2= c1 c2


c)     a1/ c2= b1/ b2= c1/ a2


Grade:12

1 Answers

jitender lakhanpal
11 years ago

Dear Pooja,

correct option is c)

solution :

it is given that the roots of the two equations are reciprocals 

n1 =  1/ m1  and   n2 = 1/ m2

m1 and m2 are the roots of

a1x2 +b1x +c1=0 

m1 + m2 = -(b1/a1)          m1*m2= c1/a1

and n1 and n2 are roots of a2x2 +b2x +c2=0  -------1

n1+n2= 1/m1+1/m2 = (m1+m2)/(m1*m2)=-(b1/c1)  ----- sum of roots 

n1*n2=  1/(m1*m2)=a1/c1-----------product of roots

now we will form the equation as x2 +(sum of roots)x +product of roots =0

x2   + (b1/c1) x +a1/c1 =0

multiply by c1 we get

c1x2   b1x +a1 =0  -------2 this is same as the equation 1 now comparing we get

we get   c1 a2  ,    b= b2    ,  a= c2

from this we get c1/a2 = 1        b1/b2= 1    a1/c2 =1   thus equating all three we get


 a1/ c2= b1/ b2= c1/ a2    so option c 


 

 

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  

jitender

askiitians expert